Documenta Mathematica

Band 15, 2010

Jianqiang Zhao Standard Relations of Multiple Polylogarithm

Values at Roots of Unity  1–34

Indranil Biswas and Norbert Hoffmann The Line Bundles on Moduli Stacks

of Principal Bundles on a Curve  35–72

José I. Burgos Gil and Răzvan Liţcanu Singular Bott-Chern Classes and the Arithmetic Grothendieck

Riemann Roch Theorem for Closed Immersions  73–176

Vanessa Miemietz and Will Turner

Hicas of Length ≤ 4  177–205

Oliver Matte On Higher Order Estimates

in Quantum Electrodynamics  207–234

Zen Harper Laplace Transform Representations and Paley–Wiener Theorems

for Functions on Vertical Strips  235–254

A. Ballester-Bolinches, J. C. Beidleman, H. Heineken and M. C. Pedraza-Aguilera Local Classes and Pairwise Mutually

Permutable Products of Finite Groups  255–265

Liangbing Jin, Yongming Liu, and Yong Zhou Blow-up of Solutions to a Periodic

Nonlinear Dispersive Rod Equation  267–283

Anna Dall’Acqua and Jan Philip Solovej Excess Charge for Pseudo-Relativistic Atoms

in Hartree-Fock Theory  285–345

Takashi Yoshinaga On the Solutions

of Quadratic Diophantine Equations  347–385

Jean-Yves Charbonnel and Anne Moreau The Index of Centralizers of

Elements of Reductive Lie Algebras  387–421

Jochen Heinloth, Alexander H. W. Schmitt The Cohomology Rings of Moduli Stacks

of Principal Bundles over Curves  423–488

Goro Shimura The Critical Values of Generalizations

of the Hurwitz Zeta Function  489–506

Jan Šťovíček Locally Well Generated

Homotopy Categories of Complexes  507–525

Feng-Yu Wang Semigroup Properties

for the Second Fundamental Form  527–543

Eike Lau Frames and Finite Group Schemes over

Complete Regular Local Rings  545–569

Adrian Vasiu and Thomas Zink Purity Results for p-Divisible Groups and Abelian Schemes over Regular Bases

of Mixed Characteristic  571–599

Miroslav Engliš, Genkai Zhang Hankel Operators and the Dixmier Trace

on Strictly Pseudoconvex Domains  601–622

Claus M. Sorensen Galois Representations Attached to

Hilbert-Siegel Modular Forms  623–670

Amir Džambić Arithmetic of a fake projective plane

and related elliptic surfaces  671–686

C. Bacuta, A. L. Mazzucato, V. Nistor, L. Zikatanov Interface and mixed boundary value problems

on n-dimensional polyhedral domains  687–745

Gerard van der Geer and Alexis Kouvidakis

The Rank-One Limit of the Fourier-Mukai Transform  747–763

Michael A. Mandell

An Inverse K-Theory Functor  765–791

Ben Moonen Special Subvarieties Arising from Families

of Cyclic Covers of the Projective Line  793–819

Jarosław Buczyński Duality and Integrability

on Contact Fano Manifolds  821–841

Yuri Prokhorov

ℚ-Fano Threefolds of Large Fano Index, I  843–872

Gerasimos Dousmanis On Reductions of Families

of Crystalline Galois Representations  873–938

Wojciech Młotkowski Fuss-Catalan Numbers

in Noncommutative Probability  939–955

Jason Starr and Johan de Jong Almost Proper GIT-Stacks

and Discriminant Avoidance  957–972

Douglas S. Bridges

Inheriting the Anti-Specker Property  973–980

Jared Weinstein Good Reduction of Affinoids

on the Lubin-Tate Tower  981–1007

Ulrich Görtz, Xuhua He Dimensions of Affine Deligne-Lusztig Varieties

in Affine Flag Varieties  1009–1028

Alejandro Adem and Zinovy Reichstein Bundles, Cohomology

and Truncated Symmetric Polynomials  1029–1047

Shun Tang A Lefschetz Fixed Point Formula for Singular Arithmetic Schemes

with Smooth Generic Fibres  1049–1108

vi